OpenStudy (anonymous):

what is 100x100(200x32)=

myininaya (myininaya):

multplication is commutative and associative we can do 100*100 first which is 10000 now we can do 10000*200=2000000 and finally we can do 2000000*32=64000000 i think i put enough zeroes there

OpenStudy (anonymous):

no i think is 16.400

OpenStudy (anonymous):

no its all multiplications only the above answer is correct

OpenStudy (anonymous):

100*100*6400 = 64,000,000